### v1

@@ Abstract
We present the first end-to-end deployment of the Gemma3 family of large language and vision models on a tiled edge dataflow architecture (AMD Ryzen AI NPU). Our work introduces a set of hardware-aware techniques. For prefill, we introduce an efficient dequantization engine, optimize tiled matrix multiplication kernels, and propose FlowQKV, a chunked, pipelined attention mechanism. For decoding, we introduce FusedDQP, which fuses dequantization and projection into a single kernel, and FlowKV, which re-structures attention to sustain high memory bandwidth utilization. Together with a compact Q4NX 4-bit quantization format, these methods yield up to $5.2\times$ faster prefill and $4.8\times$ faster decoding versus the iGPU, and $33.5\times$ and $2.2\times$ over the CPU, respectively. Power efficiency improves by as much as $67.2\times$ and $222.9\times$ compared to the iGPU and CPU. The proposed approach demonstrates that modern NPUs can deliver practical, low-power LLM and VLM inference at the edge, and provides a generalizable blueprint for mapping transformer-based models onto tiled dataflow accelerators.
@@ 1 Introduction
We evaluate 4-bit quantized Gemma3 models (1B and 4B) on an AMD Ryzen AI 7 350 NPU. Using the same platform’s iGPU and CPU as baselines, the NPU achieves, for prefill, up to $5.2\times$ over the iGPU and $33.5\times$ over the CPU. For decoding—especially at longer sequences—the NPU matches or exceeds both baselines, reaching up to $4.8\times$ over the iGPU and $2.2\times$ over the CPU. For the vision tower (processing and understanding; 4B only), the NPU is $1.7\times$ faster than the iGPU and $8.7\times$ faster than the CPU. Importantly, the NPU demonstrates markedly superior energy efficiency—achieving up to $67.2\times$ and $222.9\times$ improvements over the iGPU and CPU, respectively, as measured in Tokens Per Second per Watt (TPS/W).
@@ 4.1 Implementation
The full implementation and low-level kernel optimization are carried out using AMD’s AIE-MLIR infrastructure along with the IRON Python interface AMD/Xilinx (2025); Hunhoff et al. (2025). Link: (*The GitHub link is currently hidden to preserve anonymity during the review process.)
@@ 4.2 Evaluation and Benchmarking Results
For vision-tower inference, NPU TTFT is 4.41s vs 7.45s (iGPU) and 38.55s (CPU), yielding $1.7\times$ and $8.7\times$ speedups, respectively.
@@
To summarize, for prefill, the 1B variant shows NPU power-efficiency gains of $6.6\times$$60.0\times$ over iGPU and $22.2\times$$43.7\times$ over CPU across 1k–32k tokens; the 4B variant delivers $13.3\times$$67.2\times$ over iGPU and $68.2\times$$222.9\times$ over CPU across 1k–32k. For decoding, the 1B variant achieves $7.2\times$$20.4\times$ over iGPU and $4.5\times$$5.2\times$ over CPU across 1k–32k, while the 4B variant yields $9.3\times$$60.6\times$ over iGPU and $5.6\times$$13.2\times$ over CPU across 1k–128k.
@@ 6 Conclusion
We demonstrate the first end-to-end mapping of Gemma3 text–vision models onto a tiled edge NPU (AMD Ryzen AI NPU), showing that hardware-aware scheduling and bandwidth-centric kernels enable efficient, accurate, low-power inference without algorithmic changes. FlowQKV and FlowKV sustain high DRAM utilization for attention, FusedDQP removes projection overheads, and the compact Q4NX format reduces memory footprint while preserving fidelity. On Ryzen AI, these techniques yield multi-$\times$ speedups and up to $200\times$+ power-efficiency gains over CPU/iGPU baselines. The principles generalize to future NPUs and other dataflow accelerators, enabling scalable, real-time LLM/VLM inference at the edge.

### v2

@@ Abstract
We present the first end-to-end deployment of the Gemma3 family of large language and vision models on a tiled edge dataflow architecture (AMD Ryzen AI NPU). Our work introduces a set of hardware-aware techniques. For prefill, we introduce an efficient dequantization engine, optimize tiled matrix multiplication kernels, and propose FlowQKV, a chunked, pipelined attention mechanism. For decoding, we introduce FusedDQP, which fuses dequantization and projection into a single kernel, and FlowKV, which re-structures attention to sustain high memory bandwidth utilization. Together with a compact Q4NX 4-bit quantization format, these methods yield up to $7.5\times$ faster prefill and $5.9\times$ faster decoding versus the iGPU, and $23.7\times$ and $2.7\times$ over the CPU, respectively. Power efficiency improves by as much as $96.7\times$ and $157.7\times$ compared to the iGPU and CPU. The proposed approach demonstrates that modern NPUs can deliver practical, low-power LLM and VLM inference at the edge, and provides a generalizable blueprint for mapping transformer-based models onto tiled dataflow accelerators.
@@ 1 Introduction
We evaluate 4-bit quantized Gemma3 models (1B and 4B) on an AMD Ryzen AI 7 350 NPU. Using the same platform’s iGPU and CPU as baselines, the NPU achieves, for prefill, up to $7.5\times$ over the iGPU and $23.7\times$ over the CPU. For decoding—especially at longer sequences—the NPU matches or exceeds both baselines, reaching up to $5.9\times$ over the iGPU and $2.7\times$ over the CPU. For the vision tower (processing and understanding; 4B only), the NPU is $2.9\times$ faster than the iGPU and $14.8\times$ faster than the CPU. Importantly, the NPU demonstrates markedly superior energy efficiency—achieving up to $96.7\times$ and $157.7\times$ improvements over the iGPU and CPU, respectively, as measured in Tokens Per Second per Watt (TPS/W).
@@ 4.1 Implementation
The full implementation and low-level kernel optimization are carried out using AMD’s AIE-MLIR infrastructure along with the IRON Python interface AMD/Xilinx (2025); Hunhoff et al. (2025). Link: (https://github.com/FastFlowLM/FastFlowLM)
@@ 4.2 Evaluation and Benchmarking Results
For vision-tower inference, NPU TTFT is 2.6s vs 7.45s (iGPU) and 38.55s (CPU), yielding $2.9\times$ and $14.8\times$ speedups, respectively.
@@
To summarize, for prefill, the 1B variant shows NPU power-efficiency gains of $6.9\times$$69.7\times$ over iGPU and $22.9\times$$50.7\times$ over CPU across 1k–32k tokens; the 4B variant delivers $13.9\times$$96.7\times$ over iGPU and $70.6\times$$157.7\times$ over CPU across 1k–32k. For decoding, the 1B variant achieves $12.4\times$$24.2\times$ over iGPU and $6.2\times$$5.2\times$ over CPU across 1k–32k, while the 4B variant yields $11.9\times$$71.1\times$ over iGPU and $7.2\times$$15.8\times$ over CPU across 1k–128k.
@@ 6 Conclusion
We demonstrate the first end-to-end mapping of Gemma3 text–vision models onto a tiled edge NPU (AMD Ryzen AI NPU), showing that hardware-aware scheduling and bandwidth-centric kernels enable efficient, accurate, low-power inference without algorithmic changes. FlowQKV and FlowKV sustain high DRAM utilization for attention, FusedDQP removes projection overheads, and the compact Q4NX format reduces memory footprint while preserving fidelity. On Ryzen AI, these techniques yield multi-$\times$ speedups and up to $150\times$+ power-efficiency gains over CPU/iGPU baselines. The principles generalize to future NPUs and other dataflow accelerators, enabling scalable, real-time LLM/VLM inference at the edge.